40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 1 00:13:21 2014, Koji, Update, LSC, ALS X beat setup aligned 
    Reply  Fri May 2 13:03:30 2014, Jenne, Update, LSC, ALS Y beat setup aligned 
       Reply  Fri May 2 14:31:26 2014, Koji, Update, LSC, ALS Y beat setup aligned 
          Reply  Fri May 2 19:03:13 2014, Jenne, Update, LSC, ALS out of loop spectra ALS_OutOfLoop2_2May2014.pdf
Message ID: 9887     Entry time: Thu May 1 00:13:21 2014     Reply to this: 9904
Author: Koji 
Type: Update 
Category: LSC 
Subject: ALS X beat setup aligned 

I saw big misalignment on the GTRX camera, I went to the PSL table and aligned the beat beams.

I disconnected the RF out of the X beat PD and  connect an oscilloscope.
The beat amplitude was 15mVpp at the beginning and is 60mVpp right now.
I checked the alignment on this RF PD and the DC PD as well as the spot on the CCD.

The RF cable was connected again.

Jenne and I ran the ALS and scanned the arm cavity. We had the impression that the noise level of the ALS improved,
but I don't have correctly calibrated measurement. Let's do it tomorrow in the day time.

The Yarm beat alignment look awful. We should align this too.

ELOG V3.1.3-